OpenStudy (anonymous):

Solve: 21x - 24x2 + 6x = 9 -24x2 A. 1/2 B. 1/3 C. 3/2 D. 2/3

OpenStudy (anonymous):

B

jhonyy9 (jhonyy9):

27x=9 x=9/27 x=1/3 B is right